# Kineo: Calibration-Free Metric Motion Capture From Sparse RGB Cameras

### v1

@@ 8.2 Quantitative Evaluation
Table 1: Quantitative comparison of human pose estimation on EgoHumans and Human3.6M. Metrics include mean-per-joint position error in metric world space (W-MPJPE, in meters) and its Procrustes-Aligned version (PA-MPJPE, in meters). For each method, we indicate whether camera parameters are estimated, taken from ground truth, or omitted, where $R|\mathbf{t}]$ denotes the extrinsic camera parameters, $K$ the intrinsic camera matrix (focal length and principal point), and $\mathbf{d}$ the distortion coefficients. For each dataset, we highlight the best and second-best calibration-free methods. Note: For HSfM on Human3.6M, results were obtained through our own evaluation, while other values are reported as in the respective papers.
@@
EgoHumans [19]
UnCaliPose [56] 2.63 2.63 60.90 0.28 0.39 - - 0.33 0.44 - Est. GT GT
DUSt3R [55] - 1.15 11.00 0.61 0.74 - - 0.49 0.74 - Est. Est. -
MASt3R [25] 4.97 0.92 10.42 0.61 0.74 0.06 0.07 0.65 0.86 - Est. Est. -
HSfM [29] 2.09 0.75 9.35 0.72 0.89 0.32 0.46 0.75 0.91 - Est. Est. -
HAMSt3R [38] 2.33 0.40 10.24 0.77 - 0.06 - 0.75 - - Est. Est. -
@@
Table 2: Quantitative comparison of camera pose estimation on EgoHumans and Human3.6M. Metrics include translation error (TE, in meters), scale-aligned translation error (s-TE, in meters), angular error (AE, in degrees), Camera Center Accuracy (CCA, as a proportion), scale-aligned CCA (s-CCA, as a proportion), and Relative Rotation Accuracy (RRA, as a proportion). For each dataset we report first and second calibration-free methods. Note: For HSfM on Human3.6M, results were obtained through our own evaluation, while other values are reported as in the respective papers.
@@
Figure 9: Automatic calibration runtimes for different keypoint detectors across sequences of varying durations. The comparison includes three keypoint sources: NLF (batch size 1, full precision), RTMPose (batch size 1, full precision), and RTMPose (batch size 32, half precision). Sequence sizes are small (S, 7min10s total across 4 views), medium (M, 43min03s total across 4 views), and long (L, 1h23min total across 4 views). Each stacked bar breaks down the runtime by processing stage: Intrinsics Initialization (yellow), Extrinsics Initialization (orange), Bundle Adjustment (pink), Global Scaling (purple), Keypoints Pairs Sampling (blue), and Keypoints Detection (gray). Total runtime is indicated above each bar. RTMPose (B32 Half) consistently achieves the fastest processing times, completing all stages in less time than the total duration of the original video sequences. Importantly, the calibration runtime scales efficiently, showing only moderate increases as sequence duration grows.
@@
Kineo achieves substantial improvements in both camera estimation and 3D human reconstruction across datasets. On EgoHumans, it surpasses the strongest prior calibration-free baseline, HSfM [29]. In the fully calibration-free setting, where extrinsics, intrinsics, and distortion parameters are all estimated, Kineo achieves a camera translation error of 0.34 m (compared to 2.09 m for HSfM and 2.33 m for HAMSt3R) and an angular error of 0.69 (compared to 9.35 for HSfM and 10.24 for HAMSt3R), while achieving a vertical field-of-view error of 1.03. These improvements in camera estimation directly enhance human reconstruction quality, attaining a W-MPJPE of 0.17 m (compared to 1.04 m for HSfM and 3.80 m for HAMSt3R) and a PA-MPJPE of 2 cm (compared to 5 cm for HSfM and 14 cm for HAMSt3R). A similar trend is observed on Human3.6M, where Kineo again delivers consistent gains over previous methods. Under the same calibration-free conditions, Kineo achieves a camera translation error of 0.12 m (compared to 0.83 m for HSfM) and an angular error of 0.89 (compared to 6.44 for HSfM). The field-of-view error is similarly reduced to 0.43 (compared to 36.97 for HSfM). These improvements in camera estimation directly translate to enhanced human reconstruction quality, yielding a W-MPJPE of 0.04 m (compared to 0.47 m for HSfM) and a PA-MPJPE of 2 cm (compared to 5 cm for HSfM).
@@
To better understand the impact of estimating each camera parameter, we also examined intermediate configurations of Kineo. Providing ground-truth intrinsics, which can often be obtained from manufacturer specifications or post-capture, consistently improves both camera and human metrics, as expected. On Human3.6M, for instance, access to known intrinsics reduces translation and angular errors from 0.12 m and 0.89 to 0.01 m and 0.20, and halves the W-MPJPE from 0.04 m to 0.02 m. This performance is competitive with that of fully calibrated methods such as AdaFuse [62] and Iskakov et al. [17], indicating that Kineo attains state-of-the-art accuracy with calibrated method, even in the absence of camera extrinsics. Most notably, modeling lens distortion proves critical, especially for EgoHumans, where the wide-angle GoPro cameras introduce significant distortions. When distortion parameters are estimated, Kineo’s camera translation error decreases from 0.76 m to 0.34 m and angular error from 3.48 to 0.69, with corresponding improvements in reconstruction accuracy, W-MPJPE from 0.41 m to 0.17 m and PA-MPJPE from 0.03 m to 0.02 m. These results highlight that modeling distortions is essential for achieving high-fidelity calibration-free human reconstruction in challenging real-world capture conditions, while also providing the practical benefit of enabling accurate image undistortion for downstream tasks (see Sec. 6.1).
@@
In addition to its accuracy gains, Kineo is computationally efficient and scalable. Figure 9 reports runtimes for different pipeline stages across sequences of varying lengths and four views using multiple keypoint detectors. We evaluated three sequences of increasing length: a small sequence (S) typical of standard datasets, approximately 1 minute in duration; a medium sequence (M) of roughly 10 minutes; and a long sequence (L) of approximately 20 minutes. Three keypoint detection configurations were tested: NLF [39], RTMPose [18] with a batch size of 1 and full precision, and RTMPose with a batch size of 32 and half precision. All experiments using SMPL-based global scaling enabled and were conducted on a single NVIDIA A100 GPU. Across all configurations, keypoint detection is the most time-consuming stage, while the remaining calibration stages scale moderately with sequence length. These results indicate that Kineo’s calibration pipeline is both efficient and scalable, making it suitable for long-duration, real-world captures. For example, when processing a 20-minute, four-view recording using the RTMPose model for keypoint detection (half precision, batch size = 32), the total processing time is only $36$ minutes, substantially shorter than the cumulative duration of the input footage ($4\times 20\penalty 10000\ \text{min}=1\penalty 10000\ \text{h}\penalty 10000\ 20\penalty 10000\ \text{min}$).
8.3 Ablation Study
8.3.1 Impact of Camera Parameter Estimation
@@
Table 3: Ablation study for different combinations of ground truth (GT), estimated (est) parameters and omitted parameters, evaluated on Human3.6M and EgoHumans datasets. We isolate the effect of camera parameters estimation by using ground truth keypoints to simulate an ideal detector and ground-truth global scale.
@@
8.3.2 Impact of 2D Keypoint Detector
@@
EgoHumans [19] Ground-truth 0.007
NLF [39] 0.019
DWPose [59] 0.024
Human3.6M [16] Ground-truth 0.002
NLF [39] 0.019
RTMPose [18] 0.022
Table 4: Impact of 2D detector on the PA-MPJPE (lower is better). All camera parameters are estimated (extrinsics, intrinsics, distortion coefficients).
@@
When comparing detectors, NLF [39] consistently outperforms DWPose [59] on the EgoHumans dataset and RTMPose [18] on Human3.6M. This demonstrates that, while NLF incurs a higher computational cost, it yields lower errors in human pose estimation. This flexibility in the choice of the keypoints detector allows Kineo to balance the trade-off between accuracy and efficiency based on the specific use case: whether prioritizing maximal accuracy with a more resource-intensive model or opting for a lightweight, faster detector to achieve real-time inference of 2D keypoints.
@@
8.3.3 Impact of Global Scaling Strategy
@@
EgoHumans [19] SMPL-based 0.138 0.252
@@
Human3.6M [16] SMPL-based 0.021 0.103
@@
Table 5: Ablation study on the impact of different global scale estimation strategies (SMPL-based and metric depth-based using MoGe [54]) on reconstruction accuracy (W-MPJPE) and camera translation error (TE). All camera parameters and 2D keypoints are fixed to ground-truth values, isolating the effect of scale estimation. Results are reported for EgoHumans and Human3.6M datasets.
8.4 Qualitative Evaluation
@@
Figure 10: Qualitative comparison between HSfM (red) and Kineo (blue) on their respective best-performing sequences from Human3.6M. Ground truth camera and skeletons are shown in black. Visualization is scale-aligned.
@@
Figure 11: Qualitative results on sequences from the EgoHumans dataset. Kineo estimations are visualized in blue, while ground truth cameras and skeletons are shown in black. The scene pointmap is predicted as a downstream task of our method using MoGe [54].
Figure 11 illustrates qualitative results obtained using Kineo on representative sequences from the EgoHumans dataset, including activities such as badminton, basketball, fencing, lego assembly, tagging, tennis, and volleyball. For each sequence, the scene pointmaps were reconstructed from a single camera view. The images were first undistorted using the estimated distortion coefficients, and the estimated focal length was subsequently provided to MoGe [54] to enhance the accuracy of the reconstruction. Qualitatively, Kineo produces estimated results (in blue) that closely align with the ground truth (in black).
@@
Figure 12: Qualitative comparison on the EgoHumans basketball sequence, showing the distorted input image (left) and the undistorted image obtained using Kineo’s estimated distortion coefficients (right). Black lines denote ground-truth distorted/undistorted lines, while red lines indicate the undistorted lines reconstructed from Kineo’s estimates, which closely align with the ground truth.
9 Practical Applications
@@
9.1 In-The-Wild Capture (Offline Mode)
@@
Figure 13: Examples of craftsmen recorded for at an experimental archaeology worksite. Both sequences were captured using six uncalibrated and unsynchronized GoPro Hero11 Black cameras, with no calibration procedure necessary.
9.2 Realtime Capture (Online Mode)
@@ 10 Conclusion
Looking forward, we identify several promising research directions. First, integrating the proposed 3D confidence scores into temporal filtering could enhance robustness and temporal coherence. Second, camera pose estimation could be further refined by incorporating a photometric loss, following Choi et al. [6], who demonstrated that dynamic NeRF optimization can be used to jointly refine camera poses and reconstruct volumetric representations of the scenes.
@@ Acknowledgements.
Supplementary Materials
Sequence Original Image Undistorted Image (Ground-truth) Undistorted Image (Estimated)
Tagging [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Lego [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Fencing [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Basketball [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Volleyball [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Badminton [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Tennis [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Qualitative comparison of distorted input images (left), undistorted images using ground-truth distortion coefficients (middle), and undistorted images using estimated coefficients (right) across multiple sequences from the EgoHumans dataset. Black lines indicate ground-truth distorted/undistorted lines, while red lines show the undistorted lines obtained with distortion parameters estimated by Kineo.
{longtblr}
[ caption = Comparison of the W-MPJPE metric (lower is better) across Human3.6M sequences for Kineo (blue), HSfM (red), and ground truth (black). For each sequence, the table reports the minimum, median, maximum, mean, and standard deviation of W-MPJPE over all frames, along with visualizations of the corresponding frames. All estimates are scale-aligned for visualization. Overall, Kineo achieves lower W-MPJPE and demonstrates greater robustness with smaller standard deviation compared to HSfM., ] width=cells=valign=m,halign=c, colspec=Q[c,1.8cm] *3X[c] Q[c,1.8cm] *3X[c] Q[c,1.8cm], hline1,Z = 1pt, hline3 = 0.4pt, vline2-9 = 0.4pt, rowhead=2 Sequence HSfM W-MPJPE Kineo W-MPJPE
[…]

### v2

@@ 8.2. Quantitative Evaluation
Table 1. Quantitative comparison of human pose estimation on EgoHumans and Human3.6M. Metrics include mean-per-joint position error in metric world space (W-MPJPE, in meters) and its Procrustes-Aligned version (PA-MPJPE, in meters). For each method, we indicate whether camera parameters are estimated, taken from ground truth, or omitted, where $R|\mathbf{t}]$ denotes the extrinsic camera parameters, $K$ the intrinsic camera matrix (focal length and principal point), and $\mathbf{d}$ the distortion coefficients. For each dataset, we highlight the best and second-best calibration-free methods. Note: For HSfM on Human3.6M, results were obtained through our own evaluation, while other values are reported as in the respective papers.
@@
EgoHumans (Khirodkar et al., 2023)
UnCaliPose (Xu and Kitani, 2022) 2.63 2.63 60.90 0.28 0.39 - - 0.33 0.44 - Est. GT GT
DUSt3R (Wang et al., 2024) - 1.15 11.00 0.61 0.74 - - 0.49 0.74 - Est. Est. -
MASt3R (Leroy et al., 2024) 4.97 0.92 10.42 0.61 0.74 0.06 0.07 0.65 0.86 - Est. Est. -
HSfM (Müller et al., 2025) 2.09 0.75 9.35 0.72 0.89 0.32 0.46 0.75 0.91 - Est. Est. -
HAMSt3R (Rojas et al., 2025) 2.33 0.40 10.24 0.77 - 0.06 - 0.75 - - Est. Est. -
@@
Table 2. Quantitative comparison of camera pose estimation on EgoHumans and Human3.6M. Metrics include translation error (TE, in meters), scale-aligned translation error (s-TE, in meters), angular error (AE, in degrees), Camera Center Accuracy (CCA, as a proportion), scale-aligned CCA (s-CCA, as a proportion), and Relative Rotation Accuracy (RRA, as a proportion). For each dataset we report first and second calibration-free methods. Note: For HSfM on Human3.6M, results were obtained through our own evaluation, while other values are reported as in the respective papers.
@@
Figure 9. Automatic calibration runtimes for different keypoint detectors across sequences of varying durations. The comparison includes three keypoint sources: NLF (batch size 1, full precision), RTMPose (batch size 1, full precision), and RTMPose (batch size 32, half precision). Sequence sizes are small (S, 7min10s total across 4 views), medium (M, 43min03s total across 4 views), and long (L, 1h23min total across 4 views). Each stacked bar breaks down the runtime by processing stage: Intrinsics Initialization (yellow), Extrinsics Initialization (orange), Bundle Adjustment (pink), Global Scaling (purple), Keypoints Pairs Sampling (blue), and Keypoints Detection (gray). Total runtime is indicated above each bar. RTMPose (B32 Half) consistently achieves the fastest processing times, completing all stages in less time than the total duration of the original video sequences. Importantly, the calibration runtime scales efficiently, showing only moderate increases as sequence duration grows.
@@
Kineo achieves substantial improvements in both camera estimation and 3D human reconstruction across datasets. On EgoHumans, it surpasses the strongest prior calibration-free baseline, HSfM (Müller et al., 2025). In the fully calibration-free setting, where extrinsics, intrinsics, and distortion parameters are all estimated, Kineo achieves a camera translation error of 0.34 m (compared to 2.09 m for HSfM and 2.33 m for HAMSt3R) and an angular error of 0.69 (compared to 9.35 for HSfM and 10.24 for HAMSt3R), while achieving a vertical field-of-view error of 1.03. These improvements in camera estimation directly enhance human reconstruction quality, attaining a W-MPJPE of 0.17 m (compared to 1.04 m for HSfM and 3.80 m for HAMSt3R) and a PA-MPJPE of 2 cm (compared to 5 cm for HSfM and 14 cm for HAMSt3R). A similar trend is observed on Human3.6M, where Kineo again delivers consistent gains over previous methods. Under the same calibration-free conditions, Kineo achieves a camera translation error of 0.12 m (compared to 0.83 m for HSfM) and an angular error of 0.89 (compared to 6.44 for HSfM). The field-of-view error is similarly reduced to 0.43 (compared to 36.97 for HSfM). These improvements in camera estimation directly translate to enhanced human reconstruction quality, yielding a W-MPJPE of 0.04 m (compared to 0.47 m for HSfM) and a PA-MPJPE of 2 cm (compared to 5 cm for HSfM).
@@
To better understand the impact of estimating each camera parameter, we also examined intermediate configurations of Kineo. Providing ground-truth intrinsics, which can often be obtained from manufacturer specifications or post-capture, consistently improves both camera and human metrics, as expected. On Human3.6M, for instance, access to known intrinsics reduces translation and angular errors from 0.12 m and 0.89 to 0.01 m and 0.20, and halves the W-MPJPE from 0.04 m to 0.02 m. This performance is competitive with that of fully calibrated methods such as AdaFuse (Zhang et al., 2021) and Iskakov et al. (Iskakov et al., 2019), indicating that Kineo attains state-of-the-art accuracy with calibrated method, even in the absence of camera extrinsics. Most notably, modeling lens distortion proves critical, especially for EgoHumans, where the wide-angle GoPro cameras introduce significant distortions. When distortion parameters are estimated, Kineo’s camera translation error decreases from 0.76 m to 0.34 m and angular error from 3.48 to 0.69, with corresponding improvements in reconstruction accuracy, W-MPJPE from 0.41 m to 0.17 m and PA-MPJPE from 0.03 m to 0.02 m. These results highlight that modeling distortions is essential for achieving high-fidelity calibration-free human reconstruction in challenging real-world capture conditions, while also providing the practical benefit of enabling accurate image undistortion for downstream tasks (see Sec. 6.1).
@@
In addition to its accuracy gains, Kineo is computationally efficient and scalable. Figure 9 reports runtimes for different pipeline stages across sequences of varying lengths and four views using multiple keypoint detectors. We evaluated three sequences of increasing length: a small sequence (S) typical of standard datasets, approximately 1 minute in duration; a medium sequence (M) of roughly 10 minutes; and a long sequence (L) of approximately 20 minutes. Three keypoint detection configurations were tested: NLF (Sárándi and Pons-Moll, 2025), RTMPose (Jiang et al., 2023) with a batch size of 1 and full precision, and RTMPose with a batch size of 32 and half precision. All experiments using SMPL-based global scaling enabled and were conducted on a single NVIDIA A100 GPU. Across all configurations, keypoint detection is the most time-consuming stage, while the remaining calibration stages scale moderately with sequence length. These results indicate that Kineo’s calibration pipeline is both efficient and scalable, making it suitable for long-duration, real-world captures. For example, when processing a 20-minute, four-view recording using the RTMPose model for keypoint detection (half precision, batch size = 32), the total processing time is only $36$ minutes, substantially shorter than the cumulative duration of the input footage ($4\times 20~\text{min}=1~\text{h}~20~\text{min}$).
8.3. Ablation Study
8.3.1. Impact of Camera Parameter Estimation
@@
Table 3. Ablation study for different combinations of ground truth (GT), estimated (est) parameters and omitted parameters, evaluated on Human3.6M and EgoHumans datasets. We isolate the effect of camera parameters estimation by using ground truth keypoints to simulate an ideal detector and ground-truth global scale.
@@
8.3.2. Impact of 2D Keypoint Detector
@@
EgoHumans (Khirodkar et al., 2023) Ground-truth 0.007
NLF (Sárándi and Pons-Moll, 2025) 0.019
DWPose (Yang et al., 2023) 0.024
Human3.6M (Ionescu et al., 2014) Ground-truth 0.002
NLF (Sárándi and Pons-Moll, 2025) 0.019
RTMPose (Jiang et al., 2023) 0.022
Table 4. Impact of 2D detector on the PA-MPJPE (lower is better). All camera parameters are estimated (extrinsics, intrinsics, distortion coefficients).
@@
When comparing detectors, NLF (Sárándi and Pons-Moll, 2025) consistently outperforms DWPose (Yang et al., 2023) on the EgoHumans dataset and RTMPose (Jiang et al., 2023) on Human3.6M. This demonstrates that, while NLF incurs a higher computational cost, it yields lower errors in human pose estimation. This flexibility in the choice of the keypoints detector allows Kineo to balance the trade-off between accuracy and efficiency based on the specific use case: whether prioritizing maximal accuracy with a more resource-intensive model or opting for a lightweight, faster detector to achieve real-time inference of 2D keypoints.
@@
8.3.3. Impact of Global Scaling Strategy
@@
EgoHumans (Khirodkar et al., 2023) SMPL-based 0.138 0.252
@@
Human3.6M (Ionescu et al., 2014) SMPL-based 0.021 0.103
@@
Table 5. Ablation study on the impact of different global scale estimation strategies (SMPL-based and metric depth-based using MoGe (Wang et al., 2025b)) on reconstruction accuracy (W-MPJPE) and camera translation error (TE). All camera parameters and 2D keypoints are fixed to ground-truth values, isolating the effect of scale estimation. Results are reported for EgoHumans and Human3.6M datasets.
8.4. Qualitative Evaluation
@@
Figure 10. Qualitative comparison between HSfM (red) and Kineo (blue) on their respective best-performing sequences from Human3.6M. Ground truth camera and skeletons are shown in black. Visualization is scale-aligned.
@@
Figure 11. Qualitative results on sequences from the EgoHumans dataset. Kineo estimations are visualized in blue, while ground truth cameras and skeletons are shown in black. The scene pointmap is predicted as a downstream task of our method using MoGe (Wang et al., 2025b).
Figure 11 illustrates qualitative results obtained using Kineo on representative sequences from the EgoHumans dataset, including activities such as badminton, basketball, fencing, lego assembly, tagging, tennis, and volleyball. For each sequence, the scene pointmaps were reconstructed from a single camera view. The images were first undistorted using the estimated distortion coefficients, and the estimated focal length was subsequently provided to MoGe (Wang et al., 2025b) to enhance the accuracy of the reconstruction. Qualitatively, Kineo produces estimated results (in blue) that closely align with the ground truth (in black).
@@
Figure 12. Qualitative comparison on the EgoHumans basketball sequence, showing the distorted input image (left) and the undistorted image obtained using Kineo’s estimated distortion coefficients (right). Black lines denote ground-truth distorted/undistorted lines, while red lines indicate the undistorted lines reconstructed from Kineo’s estimates, which closely align with the ground truth.
9. Practical Applications
@@
9.1. In-The-Wild Capture (Offline Mode)
@@
Figure 13. Examples of craftsmen recorded for at an experimental archaeology worksite. Both sequences were captured using six uncalibrated and unsynchronized GoPro Hero11 Black cameras, with no calibration procedure necessary.
9.2. Realtime Capture (Online Mode)
@@ 10. Conclusion
Looking forward, we identify several promising research directions. First, integrating the proposed 3D confidence scores into temporal filtering could enhance robustness and temporal coherence. Second, camera pose estimation could be further refined by incorporating a photometric loss, following Choi et al. (Choi et al., 2025), who demonstrated that dynamic NeRF optimization can be used to jointly refine camera poses and reconstruct volumetric representations of the scenes.